OpenStudy (anonymous):

solve x^2-6x=0 by graphing or by factoring. A)x=0 B)x=-6 C)x=-6, x=0 D)x=0, x=6

OpenStudy (anonymous):

I think it would be easiest to factorise it. Try taking out an x, and then see how you get on.

jimthompson5910 (jim_thompson5910):

x^2 - 6x = 0 x(x - 6) = 0 hopefully that helps

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np

jimthompson5910 (jim_thompson5910):

tell me what you got

OpenStudy (anonymous):

c

jimthompson5910 (jim_thompson5910):

close, but no

jimthompson5910 (jim_thompson5910):

x(x - 6) = 0 x = 0 or x-6 = 0 x = 0 or x = 6

OpenStudy (anonymous):

so its d then?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

i think i kept the - sign in, that's what probably messed me up

jimthompson5910 (jim_thompson5910):

yeah it gets a lot of people, but just keep in mind that x = 6 makes x-6 =0 true since x-6 = 0 6 - 6 = 0... plug in x = 6 0 = 0 ... you can see that it's true, so x = 6 is a solution to x - 6 = 0

OpenStudy (anonymous):

i get it thanks, im gona go . night

jimthompson5910 (jim_thompson5910):

ok glad it's making sense now